At what temperature would the rms speed of chlorine-35 atoms equal the following speeds? (Note: The mass of a chlorine-35 atom is 5.81 ✕ 10−26 kg.)

(a) the escape speed from Earth, 1.12 ✕ 104 m/s

(b) the escape speed from the Moon, 2.37 ✕ 103 m/s
Answer #1

From the kinetic theory Kinetic energy of particles (½mv²) = (3/2) kT
mv² = 3kT
T = mv² / 3k ----------(1)

where v = escape velocity

m = mass of atom = 5.81 ? 10-26 kg

k = Boltzmann constant 1.38-23 J/(K.particle)

a)

from equation 1

T = mv² / 3k = 5.81 ? 10-26*[1.12 ? 104]2 / 3*1.38*10-23

T = 1.760*105 K

(b)

T = mv² / 3k = 5.81 ? 10-26*[2.37 ? 103]2 / 3*1.38*10-23

T = 7.88*103 K